ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

19. History of Mathematics

1991 MS Classification: 1 Guo, Shirong, Institute for the History of Science, Inner mongolia Normal University, Huhhot, 010022, China Yang Hui’s Influence on the Mathematical Tradition of the Ming Dynasty in China There were two important aspects of the mathematical tradition in the Song and Yuan Dynasties of China. One was represented by mathematician Yang Hui’s works from 1260-1275, and the other was represented by the works of mathematicians Qin Jiushao (1247), Li Ye (1248,1259), Zhu Shijie (1303). The first aspect was carried forward, but the second aspect was lost, in the Ming Dynasty. The fact is shown in the existing mathematical works of mathematicians, such as Wu Jing (1450), Wang Wensu (1524), Cheng Dawei (1529) and others. And Yang Hui’s mathematical tradition had been kept up in the Ming Dynasty till the western knowlodge was introduced into China at the beginning of the 17th century.